kelly had 12 dollars and spent x dollars. ami had 6 dollars and collected x dollars. the two girls then had the same amount of money. form an equation and solve it to determine the value of x.
Answer by rapaljer(4671)

Kelly = 12-x
Ami = 6 + x

12-x=6+x

Add x to each side:
12-x +x = 6+x+x
12=6 + 2x

Subtract 6 from each side:
12-6=6+2x-6
6=2x
x=3 = ANSWER!

Check:
Kelly =12 - x
Kelly = 12 - 3 = 9
Ami = 6 + x
Ami = 6 + 3 = 9

R^2